Determine the pressures of two grams (each) of gaseous H2, He, N2 and Ar, where each sample is contained in a volume of five liters at 0 degrees centigrade. Which sample has the greatest pressure under the conditions stated? Which sample has the least?
Answer #1

Ideal gas law is:
P*V = n*R*T

n is number of mol = mass / molar mass

So,
P*V = (mass/molar mass)*R*T
P= (mass*R*T) / (molar mass * V)

Mass is same for all
V is same for all
T is same for all

So, the one with smaller molar mass will have greatest pressure and the one with largest molar mass will have smallest pressure

Molar mass of H2 = 2 g/mol
Molar mass of He = 4 g/mol
Molar mass of N2 = 28 g/mol
Molar mass of Ar = 40 g/mol

So,
Greatest pressure is for H2
Smallest pressure is for Ar

Answer:
H2
Ar